the amount of oil used by a ship traveling at a uniform speed varies jointly with the distance and the square of the speed. If the ship uses 200 barrels of oil in traveling 500 miles at 32 miles per hour, determine how many barrels of oil are used when the ship travels 320 miles at 16 miles per hour

Answer by ankor@dixie-net.com(22740)

the amount of oil used by a ship traveling at a uniform speed varies jointly with the distance and the square of the speed.
k * d * r^2 = b, where
k = the constant of variation
d = dist
r = speed
b = barrels consumed
:
If the ship uses 200 barrels of oil in traveling 500 miles at 32 miles per hour,
Find the constant of variation
k * 500 * 32^2 = 200
512000k = 200
k = 200/512000
k = .000390625
:
determine how many barrels of oil are used when the ship travels 320 miles at 16 miles per hour
b = .00039625 * 320 * 16^2
b = 32 barrels of oil required